# Safety-guaranteed and Goal-oriented Semantic Sensing, Communication, and Control for Robotics

### 13 Mar 2026 v1

@@ Abstract
Wirelessly-connected robotic system empowers robots with real-time intelligence by leveraging remote computing resources for decision-making. However, the data exchange between robots and base stations often overwhelms communication links, introducing latency that undermines real-time response. To tackle this, goal-oriented semantic communication (GSC) has been introduced into wirelessly-connected robotic systems to extract and transmit only goal-relevant semantic representations, enhancing communication efficiency and task effectiveness. However, existing GSC approaches focused primarily on optimizing effectiveness metrics while overlooking safety requirements, which should be treated as the top priority in real-world robotic systems. To bridge this gap, we propose safety-guaranteed and goal-oriented semantic communication for wirelessly-connected robotic system, aiming to maximize the robotic task effectiveness subject to practical operational safety requirements. We first summarize the general safety requirements and effectiveness metrics across typical robotic tasks, including robot arm grasping, unmanned aerial vehicle (UAV)-assisted tasks, and multi-robot exploration. We then systematically analyze the unique safety and effectiveness challenges faced by wirelessly-connected robotic system in sensing, communication, and control. Based on these, we further present potential safety-guaranteed and goal-oriented sensing, communication, and control solutions. Finally, a UAV target tracking case study validates that our proposed GSC solutions can significantly improve safety rate and tracking success rate by more than 2 times and 4.5 times, respectively.
@@ I Introduction
Wirelessly-connected robotic system has emerged as a new paradigm for enabling intelligent, responsive, and decentralized robotic systems. However, existing wireless networks cannot meet the stringent timeliness requirements of bidirectional communication between robots and base stations, especially in large-scale deployments. On the one hand, the uplink must support the real-time transmission of bandwidth-intensive multi-modal sensory data, where limited wireless capacity can degrade the freshness and completeness of the sensory information, leading to delayed or suboptimal decision-making. On the other hand, the downlink transmission of control and command (C&C) demands timely transmission with high update frequency and reliability, where any packet loss or congestion might impair the robot’s safe control.
@@
To address this, goal-oriented semantic communication (GSC) has been introduced as a promising communication paradigm [15], which extracts and transmits the significance or usefulness of the raw data with respect to the data exchange goal. For uplink sensory data, recent research has expanded GSC design for different data types in various tasks [9, 12]. In [9], the scene graphs were extracted from unmanned aerial vehicle (UAV)-captured images as semantic representations of the original image, to preserve key geometric relationships while enabling accurate downstream visual question answering task. In [12], a GSC framework was proposed for point cloud-based avatar reconstruction in the Metaverse, where avatar skeleton nodes are ranked based on its importance for selective transmission. For downlink C&C data, prior work has utilized GSC to selectively transmit most useful downlink packets [13, 3], thereby reducing unnecessary communication without compromising downstream control objectives. The authors in [13] introduced the value and age of information to quantify the semantic importance and freshness of C&C packets and drop less informative packets. Another study [3] developed a deep reinforcement learning (DRL) algorithm to perform temporal selection on C&C packets, which selectively discards packets that contribute little to the downstream control task.
@@
Nevertheless, real-world robotic applications demand more than just task effectiveness, they must also meet stringent safety requirements. According to industrial robotics standards [7, 2], both the US and the EU have established formal regulations that define specific safety requirements for different types of robots operating in various scenarios. For instance, in UAV target tracking, precise tracking demands the UAV stay close to the target, but maintaining a distance of less than 3 meters significantly increases collision risk. In addition to published standards, some studies in the robotics community have incorporated safety as part of the optimization objective or explicit constraints. For example, collision probability has been widely adopted as an optimization objective alongside shortest path planning in autonomous driving applications [5], while for collaborative robots, internal joint force has been used as a key safety criteria [11] to trigger emergency stop for preventing human injury. Unfortunately, existing works in the communication community on robotic communication remain focused on optimizing traditional metrics, such as throughput [12] and mean squared error (MSE) [13], without realizing that safety is a fundamental effectiveness metric across all robotic applications that should be prioritized above all other metrics.
Moreover, another unresolved yet critical issue in both the communication and robotics communities is the lack of a unified framework that jointly considers sensing, communication, and control in a closed-loop manner. The safety and effectiveness of control rely on the quality of communication and the accuracy of sensed data; the control decisions in turn influence subsequent sensing and communication; while the type of sensed data and the requirements of the downstream control task affect the communication frequency and transmission priorities. Although recent studies [9, 12, 13, 3] have explored the pairwise of two components, they typically neglect the tight coupling among all three components. Consequently, one aspect is inevitably sacrificed or insufficiently addressed, leading to suboptimal efficiency or compromising system safety.
@@
Motivated by the above, this work systematically investigates how to enable safety-guaranteed and goal-oriented semantic sensing, communication, and control co-design in wirelessly-connected robotic system, aiming to maximize the task effectiveness while satisfying safety requirements. To the best of our knowledge, this is the first work to extend communication design beyond task effectiveness by explicitly incorporating practical safety requirements into the communication objective. Our main contributions include:
@@
  • We summarize the safety requirements and effectiveness metrics for wirelessly-connected robotic systems in representative industrial robotic tasks.
@@
  • We identify the key challenges of wirelessly-connected robotic system during sensing, communication and control. To tackle these challenges, we present novel safety-guaranteed and goal-oriented sensing, communication, and control solutions, aiming to maximizing task effectiveness while satisfying safety requirements.
@@
  • We validate our proposed GSC solutions in a UAV target tracking use case. Compared with baseline, our solutions improve the safety rate by more than 2 times and tracking success rate by more than 4.5 times, respectively.
@@
TABLE I: Commonly-used sensors in wirelessly-connected robotic system
@@
Inertial Measurement Unit (IMU) Acceleration, Angular Velocity MPU6050, Xsens MTi 100-1000 Hz Low-Medium
Magnetometer Heading HMC5883L, BNO055 10-100 Hz Low
@@
Ultrasonic Sensor Distance HC-SR04, MaxBotix 1-50 Hz Low
UWB Sensor Radio Signal Decawave DW1000 10-100 Hz Low
Wi-Fi Module Radio Signal ESP8266, Intel AX200 1-10 Hz Low
@@ II System Architecture & Use Cases
Figure 1: The illustration of the wirelessly-connected robotic system.
@@ II-A System Architecture
The wirelessly-connected robotic system consists of five key components: sensors, a sensor-to-edge link, an edge server deployed at the base station, an edge-to-robot link, and a group of robots, as shown in Fig. 1. At beginning, robot-embedded sensors capture the robot’s internal motion and state, while external sensors collect environmental information, where Table I summarizes commonly-used sensors in the system and the data types they collect. Subsequently, in the sensor-to-edge link, all the collected sensor data are wirelessly transmitted to the edge server for real-time processing. Upon receiving these data, the edge server runs real-time simulations (e.g., digital twin) or control algorithms (e.g., learning-based control) to generate optimal control decisions for different industrial robots. In the edge-to-robot link, these control decisions are translated into executable commands and transmitted back to corresponding robots for action execution. We abstract this closed-loop process into three functional stages, i.e., sensing, communication, and control, to enable systematic analysis in Sec. IV. Specifically, sensing refers to the collection of raw sensor data at the robot side and its processing at the edge server. Communication encompasses both the sensor-to-edge and edge-to-robot wireless transmissions. Control involves generating C&C data at the edge server and executing them at the robot side.
@@ II-B Use Cases
The wirelessly-connected robotic system can represent a wide range of robotic applications in practice, herein we present three typical use cases to illustrate its versatility.
@@ II-B1 Robot Arm Grasping
Robot arms can grasp irregular, slippery objects in cluttered environments using multi-modal sensing and edge processing. First, images or point clouds are sent to the edge server for object detection and segmentation with vision models like YOLO or PointNet++. After identifying the target, the edge server determines the optimal grasp pose and force using grasp synthesis models such as generative grasping CNN or Contact-GraspNet. Based on joint position and velocity data transmitted from the robot’s embedded sensors, the server then plans and returns an efficient, collision-free path. During grasp execution, tactile sensors stream feedback for real-time contact stability monitoring, ensuring a secure and reliable grasp.
@@ II-B2 UAV-assisted Tasks
UAVs are widely used for parcel delivery and target tracking in industrial settings. During operation, vision, positioning, and motion data from cameras, ultrasonic sensors, and IMUs are continuously sent to the edge server for real-time processing. Leveraging their top-down view, UAVs use multi-view image fusion to improve tracking accuracy. Motion data is also processed by adaptive flight control models at the edge server, enabling dynamic path adjustments in response to obstacles or moving targets.
@@ II-B3 Multi-robot Exploration
Multiple mobile robots often collaborate to map unexplored regions, transmitting diverse sensor data to edge server for localization, motion tracking, and environmental mapping. For localization, robots use ultra-wideband (UWB) sensors for time-of-flight measurements or Wi-Fi modules for received signal strength indicator (RSSI)-based distance estimation. For motion tracking, acceleration and velocity data captured by IMU are fused with radio-based localization data to predict trajectories and prevent collisions. For mapping, vision-equipped robots detect landmarks to construct detailed maps, while non-vision robots infer spatial layouts via analyzing the RSSI from multiple fixed RFID tags. In the edge-to-robot link, the edge server plans exploration strategies and makes coordinated decisions via methods such as cooperative multi-agent DRL (MADRL) to optimize task allocation and exploration efficiency.
@@
III Safety Requirements & Effectiveness Metrics
In this section, safety requirements and effectiveness metrics are defined to evaluate the performance of the wireless-connected robotic system in the given use cases.
III-A Safety Requirements
To ensure the safe operation of the robotic system, the following summarized requirements need to be satisfied.
III-A1 Safety Distance
It refers to the minimum separation required between robots and objects to ensure safety. This can be measured using danger field, including static danger field (SDF) and kineto-static danger field (KSDF) [8]. Moreover, it can be quantified by maintaining minimum distance guarantees, which may be fixed at all times or dynamic based on real-time relative motion and velocity between robots and objects [7].
III-A2 Safety Tracking Error
It is the minimum deviation required between the robot’s intended and actual trajectories to prevent collisions with the surrounding environment. This can be quantified by the real-time average distance between the robot’s target and actual positions [2], or the lateral offset and yaw angle error [6].
III-A3 Safety Grasping Force
It represents the required force range applied by a robotic gripper/manipulator to safely handle objects without causing damage. This involves defining upper and lower grasping force thresholds, which can be quantified by building a contact force model related to stiffness, friction and weight of objects, such as the single point with friction (SPwF) contact model [10].
III-A4 Safety Speed
It refers to the maximum operating speed that ensures safety. For robot arms, ISO/TS 15066 and ISO 10218 specify that the speed of an end effector should be limited to 250 mm/s when a human enters the workspace. For mobile robots, ISO 3691-4 specifies speed limits. When obstacles are within 500 mm laterally, the maximum speed is 1.2 m/s. If frontal clearance is also $\leq$ 500 mm, the limit is reduced to 0.7 m/s. Robots lacking personnel detection are restricted to 0.3 m/s.
III-A5 Quasi-dynamic Stability
This refers to a legged robot’s ability to maintain balance during walking, running, or jumping with payloads, thereby avoiding risks to nearby equipment and personnel. It is quantified using metrics such as the Zero Moment Point (ZMP) [1], Center of Pressure (CoP), and Force Restoration Index (FRI), which assess whether resultant torques and ground reaction forces at the contact points remain within the support polygon to prevent tipping during dynamic motion.
III-B Effectiveness Metrics
Effectiveness metrics vary by tasks and are summarized for the given use cases in Table II. For robot arm grasping, key metrics include Grasp Success Rate, Execution Time, and Mean Picks Per Hour (MPPH) [14]. In UAV-assisted tasks, two scenarios are considered: target tracking and parcel delivery. For target tracking, effectiveness is measured by Tracking Accuracy, Tracking Success Rate, and Identity F1 Score (IDF1) [4]. For parcel delivery, effectiveness metrics include Delivery Accuracy, Delivery Success Rate, and Delivery Time. Lastly, for multi-robot exploration, performance is evaluated based on Mapping Time and Mapping Error.
@@ III-B Effectiveness Metrics
Object’s attributes (e.g., type, material, size, and surface texture) Image, Point cloud $\times$ - -
Object 6D pose Point cloud $\times$ - -
@@
Obstacles (e.g., walls, roads, and barriers) Image, Point cloud $\times$
Obstacle’s 6D pose Point cloud $\times$
Number of surrounding objects Image, Point cloud $\times$
Relative distance to obstacles Image, Point cloud $\times$ $\times$
Relative speed to obstacles Video $\times$ $\times$
@@
Object deformation Point Cloud $\times$ - - - -
@@
IV Challenges & Potential Solutions
In this section, we outline the key challenges in sensing, communication, and control that must be addressed to satisfy safety requirements and enhance effectiveness metrics. Then, we present promising goal-oriented semantic research directions to address these challenges, where the challenge starts with letter “C” and the potential solution start with letter “S”.

### 26 Apr 2026 v2

# Safety-aware Goal-oriented Semantic Sensing, Communication, and Control for Robotics

@@ Abstract
Wirelessly-connected robotic systems empower robots with real-time intelligence by leveraging remote computing resources for decision-making. However, the data exchange between robots and edge servers often overwhelms communication links, introducing latency that degrades task performance. To tackle this, goal-oriented semantic communication (GSC) has been introduced for wirelessly-connected robotic systems to extract and transmit only goal-relevant semantic representations. While this improves task effectiveness, it generally overlooks practical safety requirements. Meanwhile, existing robotics research often treats safety primarily as a control-level problem, without systematically considering safety across sensing, communication, and control in a closed-loop manner. To bridge this gap, we investigate how to enable safety-aware goal-oriented semantic (SA-GS) sensing, communication, and control co-design in wirelessly-connected robotic systems, aiming to maximize the robotic task effectiveness subject to practical safety requirements. We first introduce an architecture for wirelessly-connected robotic systems and representative use cases. We then summarize general safety requirements and effectiveness metrics across the use cases. Next, we systematically analyze the unique safety and effectiveness challenges in sensing, communication, and control. Based on these, we further present potential SA-GS research directions. Finally, an Unmanned Aerial Vehicle (UAV) target tracking case study validates that one of the presented SA-GS research directions, i.e., semantic-based C&C packet execution, could significantly improve safety rate and tracking success rate by more than 2 times and 4.5 times, respectively.
@@ I Introduction
Wirelessly-connected robotic system has emerged as a new paradigm for enabling intelligent, responsive, and decentralized systems. However, existing wireless networks cannot meet the stringent timeliness requirements of bidirectional communication between robots and the edge server, especially in large-scale deployments. On the one hand, the uplink must support the real-time transmission of bandwidth-intensive multi-modal sensory data, where limited wireless capacity can degrade the information freshness and completeness, leading to delayed or suboptimal decision-making. On the other hand, the downlink transmission of control and command (C&C) signals demands timely transmission with high update frequency and reliability, where any packet loss or congestion might impair the robot’s task performance.
@@
To address this, goal-oriented semantic communication (GSC) has been introduced as a promising communication paradigm [15], which extracts and transmits the important information rather than raw data with respect to the task. For the uplink, recent research has expanded GSC design for different sensory data types and various tasks [9, 12]. To preserve key geometric relationships and enable accurate downstream visual question answering, scene graphs were extracted from Unmanned Aerial Vehicle (UAV)-captured images as semantic representations of the original image in [9]. To achieve point cloud-based avatar reconstruction in the Metaverse, avatar skeleton nodes were ranked based on their importance for selective transmission in [12]. For the downlink, prior work has utilized GSC to selectively transmit and execute the most useful C&C packets [3, 13]. The study in [3] developed a deep reinforcement learning (DRL) algorithm to perform temporal selection of C&C packets, discarding packets that contribute little to the downstream control task. The authors in [13] introduced the value and age of information to quantify the semantic importance and freshness of C&C packets and drop less informative packets.
@@
Nevertheless, real-world robotic applications demand more than just task effectiveness, they must also meet stringent safety requirements. Based on industrial robotics standards [7, 2], both the US and the EU have established formal regulations that define specific safety requirements for different types of robots operating in various scenarios. For instance, in UAV target tracking, precise tracking requires the UAV to stay close to the target, but maintaining a distance of less than 3 meters significantly increases collision risk. In addition to published standards, some studies in the robotics community have incorporated safety as part of the optimization objective or explicit constraints. For example, collision probability has been widely adopted as an optimization objective alongside shortest path planning in autonomous driving applications [5], while for collaborative robots, the internal joint force has been used as a key safety criterion [11] to trigger emergency stops to prevent human injury. However, existing works in GSC focused only on optimizing effectiveness metrics, such as throughput [12] and mean squared error (MSE) [13], without realizing that safety should be prioritized.
Moreover, existing robotics research often treats safety primarily as a control-level problem, without systematically addressing safety across sensing, communication, and control in a closed-loop manner. However, these three components are inherently coupled. The safety and effectiveness of control depend on the fidelity of sensory data and the reliability of communication. In turn, control decisions shape subsequent sensing and communication processes, while the type of sensory data and the demands of downstream control tasks jointly affect communication frequency and transmission priorities. Without considering safety in the tight coupling among these three components, at least one dimension is likely to be inadequately addressed, resulting in suboptimal efficiency and potentially compromised system safety.
@@
Motivated by the above, this work systematically investigates how to enable safety-aware goal-oriented semantic (SA-GS) sensing, communication, and control co-design in wirelessly-connected robotic systems, aiming to maximize task effectiveness while prioritizing safety. To the best of our knowledge, this is the first work to extend communication design beyond task effectiveness by explicitly incorporating practical safety requirements into the communication objective. Our main contributions include:
  • We introduce an architecture for wirelessly-connected robotic systems and representative use cases.
@@
  • We summarize the safety requirements and effectiveness metrics based on the representative use cases.
@@
  • We identify the key challenges of wirelessly-connected robotic systems for sensing, communication and control. To tackle these challenges, we present novel SA-GS future directions across these three components, aiming to maximize task effectiveness subject to practical safety requirements.
@@
  • To validate the effectiveness of the proposed SA-GS directions, we implement one representative direction, namely semantic-based C&C packet execution, in a single-UAV target-tracking task. Compared with baseline, it could improve the safety rate by more than 2 times and tracking success rate by more than 4.5 times, respectively.
@@
TABLE I: Commonly-used sensors in wirelessly-connected robotic systems.
@@
Inertial Measurement Unit (IMU) Acceleration, Angular Velocity MPU6050 IMU 100-1000 Hz Low-Medium
Magnetometer Heading BNO055 Orientation Sensor 10-100 Hz Low
@@
Ultrasonic Sensor Distance MaxBotix Ultrasonic Estimator 1-50 Hz Low
UWB Sensor Radio Signal DW1000 UWB transceiver 10-100 Hz Low
Wi-Fi Module Radio Signal Intel AX200 WiFi adapter 1-10 Hz Low
@@ II System Architecture & Use Cases
Figure 1: Illustration of a wirelessly-connected robotic system.
@@ II-A System Architecture
The wirelessly-connected robotic system consists of five key components: sensors, a sensor-to-edge link, an edge server deployed at the base station, an edge-to-robot link, and a group of robots, as shown in Fig. 1. At the beginning, robot-embedded sensors capture the robot’s internal motion and state, while external sensors collect environmental information, where Table I summarizes commonly-used sensors in the system and the data types they collect. Subsequently, in the sensor-to-edge link, all the collected sensor data are wirelessly transmitted to the edge server for real-time processing. Upon receiving these data, the edge server runs real-time simulations (e.g., digital twin) or control algorithms (e.g., learning-based control) to generate optimal control decisions for different industrial robots. In the edge-to-robot link, these control decisions are translated into executable commands and transmitted back to corresponding robots for action execution. To provide a unified basis for the subsequent analysis in Sec. III and IV, we abstract this closed-loop process into three functional stages, namely sensing, communication, and control. Specifically, sensing refers to the collection of raw sensor data at the robot side and its processing at the edge server. Communication encompasses both sensor-to-edge and edge-to-robot wireless transmissions. Control involves generating C&C packets at the edge server and executing them at the robot side.
@@ II-B Use Cases
The considered wirelessly-connected robotic system can represent a wide range of robotic applications in practice. Herein, we present three typical use cases to illustrate its versatility.
@@ II-B1 Robot Arm Grasping
Robot arms can grasp irregular, slippery objects in cluttered environments using multi-modal sensing and edge processing. First, images or point clouds are sent to the edge server for object detection and segmentation with vision models like YOLO or PointNet++. After identifying the target, the edge server determines the optimal grasp pose and force using grasp synthesis models such as generative grasping CNN or Contact-GraspNet. Based on joint position and velocity data transmitted from the robot’s embedded sensors, the server then plans and returns an efficient, collision-free path. During grasp execution, tactile sensors stream feedback for real-time contact stability monitoring, ensuring a secure and reliable grasp. However, perception errors or delayed tactile feedback may cause excessive grasping force, posing safety risks to both the object and nearby operators.
@@ II-B2 UAV-assisted Tasks
UAVs are widely used for parcel delivery and target tracking in industrial settings. During operation, vision, positioning, and motion data from cameras, ultrasonic sensors, and IMUs are continuously sent to the edge server for real-time processing. Leveraging their top-down view, UAVs use multi-view image fusion to improve tracking accuracy. Motion data is also processed by adaptive flight control models at the edge server, enabling dynamic path adjustments in response to obstacles or moving targets. However, asynchronous multi-modal data fusion might prevent the UAV from maintaining a safe operating speed, thereby increasing the risk of collisions.
@@ II-B3 Multi-robot Exploration
Multiple mobile robots often collaborate to map unexplored regions, transmitting diverse sensor data to the edge server for localization, motion tracking, and environmental mapping. For localization, robots can use ultra-wideband (UWB) sensors for time-of-flight measurements or WiFi modules for received signal strength indicator (RSSI)-based distance estimation. For motion tracking, acceleration and velocity data captured by an IMU can be fused with radio-based localization data to predict trajectories and prevent collisions. For mapping, vision-equipped robots can detect landmarks to construct detailed maps, while non-vision robots infer spatial layouts via analyzing the RSSI from multiple fixed RFID tags. In the edge-to-robot link, the edge server can plan exploration strategies and make coordinated decisions via methods such as cooperative multi-agent DRL (MADRL) to optimize task allocation and exploration efficiency. However, delayed localization updates or uncoordinated decision-making may reduce safe inter-robot distance and lead to collisions in dense environments.
Although the above use cases share the same architecture, they differ substantially in sensing modalities, task objectives, and decision loops. Thus, the associated safety requirements and effectiveness metrics are also task-dependent, as detailed in Sec. III.
@@
Object’s attributes (e.g., type, material, size, and surface texture) Image, Point cloud $\mathrm{o}$ - -
Object 6D pose Point cloud $\mathrm{o}$ - -
@@
Obstacles (e.g., walls, roads, and barriers) Image, Point cloud $\mathrm{o}$
Obstacle’s 6D pose Point cloud $\mathrm{o}$
Number of surrounding objects Image, Point cloud $\mathrm{o}$
Relative distance to obstacles Image, Point cloud $\mathrm{o}$ $\mathrm{o}$
Relative speed to obstacles Video $\mathrm{o}$ $\mathrm{o}$
@@
Object deformation Point Cloud $\mathrm{o}$ - - - -
@@
III Safety Requirements & Effectiveness Metrics
In this section, safety requirements and effectiveness metrics are defined to evaluate the performance of wireless-connected robotic systems for the given use cases.
III-A Safety Requirements
To ensure the safe operation of a robotic system, the following requirements need to be satisfied.
III-A1 Safety Distance
This refers to the minimum separation required between robots and objects to ensure safety. The safety distance can be measured using the concept of danger field, including static danger field (SDF) and kineto-static danger field (KSDF) [8]. Moreover, it can be quantified by maintaining minimum distance guarantees, which may be fixed at all times or dynamic based on the real-time relative motion and velocity between robots and objects [7].
III-A2 Safety Tracking Error
This is the minimum deviation required between the robot’s intended and actual trajectories to prevent collisions with the surrounding environment. It can be quantified by the real-time average distance between the robot’s target and actual positions [2], or the lateral offset and yaw angle error [6].
III-A3 Safety Grasping Force
This represents the required force range applied by a robotic gripper/manipulator to safely handle objects without causing damage. It involves defining upper and lower grasping force thresholds, which can be quantified by building a contact force model related to the stiffness, friction, and weight of objects, such as the single point with friction (SPwF) contact model [10].
III-A4 Safety Speed
This refers to the maximum operating speed that ensures safety. For robot arms, ISO/TS 15066 and ISO 10218 specify that the speed of an end effector should be limited to 250 mm/s when a human enters the workspace. For mobile robots, ISO 3691-4 specifies speed limits. When obstacles are within 500 mm laterally, the maximum speed is 1.2 m/s. If additionally frontal clearance is $\leq$ 500 mm, the limit is reduced to 0.7 m/s. Robots lacking personnel detection are restricted to 0.3 m/s.
III-A5 Quasi-dynamic Stability
This refers to a legged robot’s ability to maintain balance during walking, running, or jumping with payloads, thereby avoiding risks to nearby equipment and personnel. It is quantified using metrics such as the Zero Moment Point (ZMP) [1], Center of Pressure (CoP), and Force Restoration Index (FRI), which assess whether the torques and ground reaction forces experienced at the contact points remain within the support polygon to prevent tipping during dynamic motion.
III-B Effectiveness Metrics
Effectiveness metrics vary by task and are summarized for the given use cases in Table II. For robot arm grasping, key metrics include Grasp Success Rate, Execution Time, and Mean Picks Per Hour (MPPH) [14]. For UAV-assisted tasks, two scenarios are considered: target tracking and parcel delivery. For target tracking, effectiveness is measured by Tracking Accuracy, Tracking Success Rate, and Identity F1 Score (IDF1) [4]. For parcel delivery, effectiveness metrics include Delivery Accuracy, Delivery Success Rate, and Delivery Time. Lastly, for multi-robot exploration, performance is evaluated based on Mapping Time and Mapping Error.
IV Challenges & Future Research Directions
In this section, we examine the key challenges in sensing, communication, and control from the perspective of improving task effectiveness while prioritizing safety. The main difficulty is that, under limited sensing, communication, and computation resources, robotic systems cannot maintain all information with equal fidelity, timeliness, and reliability. This gives rise to a fundamental trade-off: safety-critical information should be prioritized in sensing and transmission, and safety-critical commands should be executed with precedence, while the remaining resources can be used to enhance task effectiveness. Motivated by this observation, we discuss the major challenges across these three components and outline corresponding potential SA-GS research directions, which could be validated in future work. In the following discussion, we denote the key challenges by “C” and the potential research directions by “D” for clarity.